OpenStudy (anonymous):

6. Solve the equation for x: 19 = (½)x + 3 A. 40 B. 32 C. 36 D. 44

OpenStudy (gudden):

on multiplying both sides of the equation with 2 we get (19)2 = x + (3)2 38 = x + 6 taking +6 to the Right Hand side we get 38 - 6 = x = 32

OpenStudy (gudden):

so the answer is B.. so is it clear ??